# Diffusion-Based Impedance Learning for Contact-Rich Manipulation Tasks

### v2

@@ II-B Energy-based Impedance Control
Figure 2: Norton equivalent network of Impedance Control. The interaction is represented as the combination of a motion source and an impedance source. The displacement $\Delta\bm{x}$ is mapped through stiffness $\bm{K}$ to generate wrench, while the environment contributes external wrenches. This energy-based view ensures consistency between commanded motion and force response.
II-C Diffusion Model for reconstruction of sZFT
@@
From the perspective of diffusion, the displacement $\Delta\bm{x}$ can be interpreted as noise corrupting the equilibrium pose $(\bm{p}_{0},\mathbf{q}_{0})$. The role of the Diffusion Model is to reconstruct the equilibrium pose by iteratively removing this noise. During training, the model learns to predict the difference between the observed pose $(\bm{p},\mathbf{q})$ and the ground truth equilibrium pose, guided by the measured external wrench $\bm{F}_{\text{ext}}$ (Figure 3). At inference, the model denoises $\Delta\bm{x}$ and outputs the reconstructed sZFT $(\hat{\bm{p}}_{0},\hat{\mathbf{q}}_{0})$.
@@
Figure 3: Diffusion-based reconstruction of the simulated Zero-Force Trajectory (sZFT). The observed pose $\{\bm{p},\mathbf{q}\}$ together with the external wrench $\bm{F}_{\text{ext}}$ is treated as noise relative to the equilibrium $\{\bm{p}_{0},\mathbf{q}_{0}\}$. The Diffusion Model iteratively denoises this input to reconstruct the equilibrium trajectory, which is then used in energy-based stiffness estimation.
@@
It is important to distinguish between ZFT $(\bm{p}_{0},\mathbf{q}_{0})$ and sZFT $(\hat{\bm{p}}_{0},\hat{\mathbf{q}}_{0})$ in our formulation. The ZFT is the nominal Zero-Force Trajectory: a programmed equilibrium motion that remains unchanged in execution and can be generated by frame-based programming, imitation learning (for free-space motion [49, 50], with possible post-processing for contact tasks [51, 52]), or telemanipulation [53]. For deployment in the experiments of this paper, free-space trajectories were generated programmatically by chaining waypoints with minimum-jerk and Spherical Linear Interpolation (SLERP) [54], without any contact feedback. The code for ZFT generation is available on GitHub [^6] https://github.com/StrokeAIRobotics/DiffusionBasedImpedanceLearning.
The sZFT, in contrast, is not programmed. It is reconstructed by the Diffusion Model from the observed pose and external wrench and serves exclusively as input to stiffness estimation. Thus, while the ZFT prescribes the nominal equilibrium motion in the absence of contact, the sZFT provides a denoised, contact-consistent equilibrium estimate for real-time impedance adaptation. The controller continues to track the ZFT as the nominal reference, incorporating the sZFT implicitly through stiffness modulation (Figure 4).
@@
For example, when the robot encounters an unexpected obstacle such as a bumper (Figure 5, right), the planned ZFT drives it into contact, producing an external wrench. Conditioned on this wrench, the Diffusion Model reconstructs the sZFT: a shifted trajectory that is physically consistent with the contact. This reconstruction then forms the basis for updating stiffness and damping, enabling the robot to preserve the intended impedance behavior even as its motion adapts to the new environment.
@@
$\mathbf{q}_{\text{noise}}=\mathbf{q}\,\mathbf{q}_{0}^{-1},$ (11)
which encodes the rotational displacement between the observed quaternion $\mathbf{q}$ and the equilibrium quaternion $\mathbf{q}_{0}$. Partial training noise is injected to $\mathbf{q}_{0}$ using SLERP [54](Section III-B), and the network is trained to predict $\hat{\mathbf{q}}_{\text{noise}}$. At inference, the equilibrium quaternion is reconstructed by
$\hat{\mathbf{q}}_{0}=\mathbf{q}\,\hat{\mathbf{q}}_{\text{noise}}^{-1}.$ (12)
The reconstructed sZFT provides the essential input for energy-based stiffness estimation, which is the ratio between equilibrium displacement $\Delta\bm{x}$ and measured wrench $\bm{F}_{\text{ext}}$ (eq.(7)). This enables stiffness modulation to be computed in the energy domain (Section III-D). In this sense, diffusion serves as the bridge between contact-rich (“noisy”) observed trajectories and impedance-based adaptation (Figure 4).
Figure 4: Summary of Diffusion-Based Impedance Learning. Teleoperated demonstrations provide ground truth data for model training, which become corrupted by noise during interaction. The Diffusion Model reconstructs the sZFT, which then, together with the external wrench $\bm{F}_{\text{ext}}$, serves as input for energy-based stiffness estimation and adaptive Impedance Control.
III Diffusion-Based Impedance Learning
III-A Data Collection
All training data for Diffusion-Based Impedance Learning was collected through teleoperation with AVP, which provided ground truth sZFTs alongside measured wrenches and robot motion. Unlike demonstrations with manipulandums or other telemanipulation devices that are spatially restricted, this setup allowed unconstrained demonstrations. In practice, the operator’s hand pose tracked by AVP could move freely in space, unconstrained by physical objects, whereas the robot end-effector was constrained by actual contact with those same surfaces. This set-up enabled to capture both free space motion and contact-induced deviations, which is important for applications that include transitions into and out of contact.
Two datasets were generated. The first consisted of a parkour-style scenario with multiple obstacles, including a bumper, circular target, ramp, and flat tabletop (Figure 5, right). A total of 17,457 samples were recorded. These data captured diverse contact interactions representative of unstructured environments.
The second dataset focused on robotic-assisted upper-limb rehabilitation tasks (Figure 5, left). A total of 55,838 samples were recorded. Eight healthy adults (4 female, 4 male; age 20–64 years; height 162–192 cm; weight 64–91 kg) performed movements derived from standard post-stroke protocols. Each task was executed under three conditions: (i) passive, where the robot guided the subject’s arm; (ii) cooperative, where subjects actively minimized interaction wrenches; and (iii) mixed, where both robot and subject contributed. The study was approved by MIT’s Committee on the Use of Humans as Experimental Subjects (COUHES), with informed consent obtained from all participants. This dataset was used to train the model presented in this paper; however, the deployment of upper-limb rehabilitation applications is part of a separate project and is not reported here.
@@ III-B Noise Scheduling
For the denoising of the Diffusion Model, the maximum noise magnitude was computed separately for translations and rotations, and a variational noise schedule was applied across diffusion steps.
@@ III-B1 Translations
For translations, the deviation between the observed end-effector position $\bm{p}$ and the ground-truth ZFT $\bm{p}_{0}$ was treated as noise. Additional zero-mean Gaussian noise $\bm{\epsilon}_{\text{gauss}}\sim\mathcal{N}(0,\sigma^{2}I)$ was injected during training to improve generalization and prevent overfitting [55]. The translational noise for the model is defined as
@@ III-B2 Rotations
In our robot controller, orientations are represented with unit quaternions $\mathbf{q}$. Injecting noise, as in the translational case, would violate the unit-norm constraint and distort the geometry of rotations. To address this, we introduce a SLERP-based [54] noise scheduler that perturbs quaternions directly on the unit sphere. This approach preserves valid orientations, provides perturbations consistent with geodesics, and allows rotational noise to be scheduled smoothly across diffusion steps. To our knowledge, this is the first physically-consistent noise scheduler for rotations in robotics, formulated directly on a global representation of rotations [47].
The SLERP operation is illustrated in Figure 6. The interpolation between two quaternions $\mathbf{q}_{a}$ and $\mathbf{q}_{b}$ is defined as
$\text{SLERP}(\mathbf{q}_{a},\mathbf{q}_{b};\mu)=\frac{\sin((1-\mu)\Omega)}{\sin(\Omega)}\mathbf{q}_{a}+\frac{\sin(\mu\Omega)}{\sin(\Omega)}\mathbf{q}_{b},$ (15)
where $\Omega=\arccos(\langle\mathbf{q}_{a},\mathbf{q}_{b}\rangle)$ is the angle between $\mathbf{q}_{a}$ and $\mathbf{q}_{b}$, and $\mu\in[0,1]$ is the interpolation factor.
@@
III-C Model Selection to reconstruct sZFT
Our model selection was guided by two initial requirements: (i) sufficient accuracy in reconstructing both translational and rotational components and (ii) a small model capacity, suitable for real-time deployment. To evaluate this trade-off, we varied hidden dimensions, number of layers, attention heads, and loss-weighting strategies. Performance was assessed using translational error (positional loss in $m$) and rotational error (angular magnitude $\theta$ about unit-axis and unit-axis deviation $\alpha$ in degrees).
@@ III-C Model Selection to reconstruct sZFT
Our initial requirements were 2 mm for the translational error and $2.8^{\circ}$ (half of the small-angle approximation) for the angular error and unit-axis deviation. These values were determined heuristically based on other contact-rich experiments [56].
@@
Hidden Dim Heads Layers $\bm{\theta}$ Loss [deg] $\bm{\alpha}$ Loss [deg] Positional Loss [$m$]
@@
A notable result was that retraining this model on a combined dataset of parkour and robotic-assisted therapy motions further improved accuracy (positional error $0.883$ mm, $\theta$ error $0.233^{\circ}$, $\alpha$ error $0.002^{\circ}$). Importantly, the two datasets were collected with different tools and inertial properties (lightweight tool for parkour and a 600 g prosthetic hand for physical therapy) yet the model compensated for these differences without explicit training. This robustness highlights the ability of Diffusion-Based Impedance Learning to generalize across diverse physical environments. The selected architecture was ultimately adopted as the final model because it satisfied the accuracy requirements while remaining small enough for real-time torque control.
Another notable finding of this work is the small amount of training data required. The model already showed strong performance on the parkour dataset with only tens of thousands of samples. In our case, 17,45717{,}457 samples at a controller rate of 0.005s0.005\ $\text${s} correspond to just 1.5 hours of demonstration. This is in contrast to image-based Diffusion Models, which often require hundreds to thousands of demonstrations to achieve high-quality results [53, 57].
III-D Energy-based Stiffness Estimation
Based on the sZFT, we compute stiffness by equating virtual elastic energy stored in a virtual spring to measured work performed by external wrenches.
For deployment, a user can specify baseline translational and rotational stiffness values. The Diffusion Model predicts the sZFT, i.e., the equilibrium pose consistent with the current interaction wrench. Rather than altering the programmed ZFT, we preserve the intended interactive behavior by computing nonnegative reduction terms $k_{t,i}^{\star}$ and $k_{r,i}^{\star}$, which are subtracted from the predefined baseline stiffnesses $K_{t,i,\text{max}}$ and $K_{r,i,\text{max}}$, thereby keeping the effective stiffness relative to the sZFT constant:
@@ III-D Energy-based Stiffness Estimation
Here, $i\in\{x,y,z\}$ denotes the translational or rotational axis. In our experiments, the baseline stiffness values were set deliberately high ($K_{t,i,\text{max}}=800\ \text{N/m}$, $K_{r,i,\text{max}}=150\ \text{Nm/rad}$) to ensure that the robot had sufficient energy for all tasks, while allowing the model to autonomously reduce stiffness when contact occurred.
III-D1 Translations
For translations, the displacement is defined as $\bm{e}_{t}=\bm{p}-\bm{p}_{0}$. The elastic energy $E_{t}\in\mathbb{R}$ of a virtual translational spring is
$E_{t}^{\text{spring}}=\tfrac{1}{2}\bm{K}_{t}\lVert\bm{e}_{t}\rVert^{2}.$ (17)
Here, $\lVert.\lVert$ denotes the Euclidean norm. The work done by the external force $\bm{f}_{\text{ext}}\in\mathbb{R}^{3}$ along the displacement $\bm{e}_{t}$ is given by the inner product $\langle.,.\rangle$:
$E_{t}^{\text{work}}=\langle\bm{f}_{\text{ext}},\bm{e}_{t}\rangle.$ (18)
Equating eq. (17) with eq. (18) yields the translational stiffness contribution
$\bm{k}_{t}^{\star}=\frac{2\langle\bm{f}_{\text{ext}},\bm{e}_{t}\rangle}{\lVert\bm{e}_{t}\rVert^{2}}.$ (19)
To avoid oscillatory behavior, a damping term proportional to velocity error is introduced:
$\tilde{\bm{e}}_{t}=\kappa_{t}\ \bm{e}_{t}-\gamma_{t}\ \dot{\bm{e}}_{t},$ (20)
where $\kappa_{t}$ is a unit translational stiffness and $\gamma_{t}$ a damping coefficient (Appendix A-C).
III-D2 Rotations
For rotations, the unit axis–angle representation $\bm{e}_{r}=\bm{u}_{0}\ \theta_{0}$ is used (eq. (5)). The corresponding elastic energy $E_{r}\in\mathbb{R}$ of the virtual rotational spring is
$E_{r}^{\text{spring}}=\tfrac{1}{2}\bm{K}_{r}\lVert\bm{e}_{r}\rVert^{2}.$ (21)
The work done by the external moment $\bm{m}_{\text{ext}}\in\mathbb{R}^{3}$ about the rotational displacement $\bm{e}_{r}$ can be calculated by
$E_{r}^{\text{work}}=\langle\bm{m}_{\text{ext}},e_{r}\rangle.$ (22)
Equating eq. (21) with eq. (22) results in the rotational stiffness contribution
$\bm{k}_{r}^{\star}=\frac{2\langle\bm{m}_{\text{ext}},\bm{e}_{r}\rangle}{\lVert\bm{e}_{r}\rVert^{2}}.$ (23)
Dynamic effects are addressed by augmenting the error with a damping term:
$\tilde{\bm{e}}_{r}=\kappa_{r}\ \bm{e}_{r}-\gamma_{r}\ \bm{\omega},$ (24)
where $\kappa_{r}$ is a unit rotational stiffness, $\bm{\omega}$ is the angular velocity, and $\gamma_{r}$ a damping coefficient (Appendix A-C).
III-D3 Algorithm

### v3

@@ II-B3 Control Moment
Figure 2: Norton equivalent network to represent physical interaction as the combination of a virtual motion source and an impedance source. The displacement $\Delta\bm{x}$ is coupled to stiffness $\bm{K}$ to generate wrench, while the environment contributes external wrenches.
II-B4 Control Torque
Finally, the transpose of the Jacobian $\bm{J}(\bm{q}):\mathbb{R}^{n}\rightarrow\mathfrak{se}(3)$ provides the mapping between the 6-dimensional wrenches and the commanded joint torques $\bm{\tau}_{\text{ctrl}}\in\mathbb{R}^{n}$:
$\bm{\tau}_{\text{ctrl}}=\bm{J}(\bm{q})^{T}\begin{pmatrix}\bm{f}_{\text{ctrl}}\\[2.0pt] \bm{m}_{\text{ctrl}}\end{pmatrix}$. (15)
In addition to the control wrench, the environment contributes external wrenches $\bm{F}_{\text{ext}}\in\mathfrak{se}^{\ast}(3)$.
The controller implemented in this paper can be found in our Github repository.
II-B5 Norton Equivalent Network
The Impedance Control framework can be illustrated by the Norton Equivalent Network [21, 38], as represented in Figure 2. Here, two sources interact to manage physical interaction:
  • The motion source, which represents a sequence of equilibrium poses $\{\bm{p}_{0},\bm{Q}_{0}\}$ that together form the ZFT.
  • The impedance source, which represents the baseline stiffness $\bm{K}$.
The motion source is part of the information domain. For contact-rich manipulation, it is a virtual nominal trajectory that lies beyond the physical constraints imposed by the environment. Impedance couples the motion source in the information domain to the energy domain (energy exchange through $\bm{F}$ and $\{\dot{\bm{p}},\bm{\omega}\}$), a key property that we exploit later for energy-based stiffness estimation (Section III-E).
III Diffusion-Based Impedance Learning
III-A Diffusion Model for Reconstruction of sZFT
@@
From the perspective of diffusion, the displacement $\Delta\bm{x}$ can be interpreted as noise corrupting the equilibrium pose $(\bm{p}_{0},\bm{Q}_{0})$. The role of the Diffusion Model is to reconstruct the equilibrium pose by iteratively removing this noise. During training, the model learns to predict the difference between the observed pose $(\bm{p},\bm{Q})$ and the ground truth equilibrium pose, guided by the measured external wrench $\bm{F}_{\text{ext}}$ (Figure 3). At inference, the model denoises $\Delta\bm{x}$ and outputs the reconstructed sZFT $(\hat{\bm{p}}_{0},\hat{\bm{Q}}_{0})$.
@@
Figure 3: Diffusion-based reconstruction of the simulated Zero-Force Trajectory (sZFT). The observed pose $\{\bm{p},\bm{Q}\}$ together with the external wrench $\bm{F}_{\text{ext}}$ is treated as noise relative to the equilibrium $\{\bm{p}_{0},\bm{Q}_{0}\}$. The Diffusion Model iteratively denoises this input to reconstruct the equilibrium trajectory, which is then used in energy-based stiffness estimation.
Throughout this paper, all quantities denoted by a hat (^)($\hat{\cdot}\,$) are predictions of the Diffusion Model and correspond to the sZFT.
@@
It is important to distinguish between ZFT $(\bm{p}_{0},\bm{Q}_{0})$ and sZFT $(\hat{\bm{p}}_{0},\hat{\bm{Q}}_{0})$ in our formulation. The ZFT is the nominal Zero-Force Trajectory: a programmed nominal trajectory that remains unchanged in execution and can be generated by frame-based programming, imitation learning (for free-space motion [57, 26], with possible post-processing for contact tasks [50, 49]), or teleoperation [10]. For deployment in the experiments of this paper, ZFTs were generated programmatically by chaining waypoints with minimum-jerk and Spherical Linear Interpolation (SLERP) [59]. A video of the demonstrations is available on the project website [^6] https://diffusion-imp-anon.GitHub.io/DiffusionBasedImpedanceLearning/ and the code for ZFT generation is available on GitHub [^7] https://GitHub.com/diffusion-imp-anon/DiffusionBasedImpedanceLearning/tree/main/Deployment/TrajectoryGeneration.
The sZFT, on the other hand, provides a denoised, contact-consistent equilibrium estimate. It forms the basis for updating stiffness and damping (Section III-E), enabling the robot to preserve the intended interactive behavior even for unknown contact with the environment. During deployment, the controller continues to track the ZFT as the nominal reference and incorporates stiffness modulation implicitly through the change of the reconstructed sZFT.
@@
For example, when the robot encounters an unexpected obstacle, the planned ZFT drives the end-effector into contact, producing an external wrench. Conditioned on this wrench, the Diffusion Model reconstructs the sZFT and updates the controlled impedance. An illustration of the difference between ZFT and sZFT can be seen in Figure 4.
Figure 4: Illustration of the nominal Zero-Force Trajectory during deployment (ZFT, light blue) and the reconstructed simulated Zero-Force Trajectory (sZFT, skin color), generated by the Diffusion Model. The ground truth sZFT was created by teleoperation, which tracked the wrist frame {W} and serves as an input to the robot controller to move the end-effector frame {EE}. In the illustration, the ZFT is just a straight line. However, due to the external wrench during contact, the Diffusion Model will reconstruct the sZFT to imitate the demonstrated behavior. Autonomously, the robot will adapt its compliant behavior through impedance to smoothly go over the obstacle.
@@
$\bm{Q}_{\text{noise}}=\bm{Q}\,\bm{Q}_{0}^{-1},$ (18)
which encodes the rotational displacement between the observed quaternion $\bm{Q}$ and the equilibrium quaternion $\bm{Q}_{0}$. Partial training noise is injected to $\bm{Q}_{0}$ using SLERP [59](Section III-B), and the network is trained to predict $\hat{\bm{Q}}_{\text{noise}}$. At inference, the equilibrium quaternion is reconstructed by
$\hat{\bm{Q}}_{0}=\bm{Q}\,\hat{\bm{Q}}_{\text{noise}}^{-1}.$ (19)
The reconstructed sZFT provides the key link between the Diffusion Model and impedance adaptation. It serves as the input for the energy-based stiffness estimation described in Section III-E.
@@ III-B Noise Scheduling
For the denoising process, the maximum noise magnitude was computed separately for translations and rotations, and a variational noise schedule was applied across diffusion steps.
@@ III-B1 Translations
For translations, the deviation between the observed end-effector position $\bm{p}$ and the ground-truth ZFT $\bm{p}_{0}$ was treated as noise. Additional zero-mean Gaussian noise $\bm{\epsilon}_{\text{gauss}}\sim\mathcal{N}(0,\sigma^{2}I)$ was injected during training to improve generalization and prevent overfitting [6]. The translational noise for the model was defined as
@@ III-B2 Rotations
In our robot controller, orientations are represented with unit quaternions $\bm{Q}$. Injecting noise, as in the translational case, would violate the unit-norm constraint and distort the geometry of rotations. To address this, we introduce a SLERP-based [59] noise scheduler that perturbs quaternions directly on the unit sphere. This approach preserves valid orientations, provides perturbations consistent with geodesics, and allows rotational noise to be scheduled smoothly across diffusion steps. To our best knowledge, this is the first physically-consistent noise scheduler for rotations in robotics, formulated directly on a global representation of rotations [38].
The SLERP process is illustrated in Figure 5. The interpolation between two quaternions $\bm{Q}_{a}$ and $\bm{Q}_{b}$ is defined as
$\text{SLERP}(\bm{Q}_{a},\bm{Q}_{b};\mu)$ $=\frac{\sin((1-\mu)\Omega)}{\sin(\Omega)}\bm{Q}_{a}$ (22)
$\quad+\frac{\sin(\mu\Omega)}{\sin(\Omega)}\bm{Q}_{b}.$
where $\Omega=\arccos(\langle\bm{Q}_{a},\bm{Q}_{b}\rangle)$ is the angle between $\bm{Q}_{a}$ and $\bm{Q}_{b}$, and $\mu\in[0,1]$ is the interpolation factor.
@@
III-C Data Collection
All training data for Diffusion-Based Impedance Learning were collected through teleoperation using AVP, which provided ground-truth sZFTs together with measured interaction wrenches and recorded end-effector poses. The teleoperation framework enabled free-space motion, contact scenarios, and transitions into and out of contact. During data collection, AVP tracked the operator’s wrist pose, which was used as the commanded ZFT during teleoperation and subsequently treated as the ground-truth sZFT for model training. In parallel, external wrenches at the robot end-effector were recorded using a force/torque sensor.
Unlike demonstrations with manipulandums and haptic interfaces, the AVP framework allowed spatially unrestricted teleoperation. In practice, the operator’s hand pose tracked by AVP could move freely in space, unconstrained by physical objects, whereas the robot end-effector was constrained by the actual contact surfaces. Therefore, the shift between the commanded ZFT and the robot pose generated an external wrench.
Two datasets were generated. The first consisted of a parkour-style scenario with multiple obstacles (Figure 6, right). A total of 17,457 samples were recorded. These data captured diverse interaction behaviors, including free-space motion, transitions into and out of contact, and sustained contact with unknown obstacles in an unstructured environment.
The second dataset focused on robotic-assisted upper-limb rehabilitation tasks (Figure 6, left). A total of 55,838 samples were recorded. Eight healthy adults (4 female, 4 male; age 20–64 years; height 162–192 cm; weight 64–91 kg) performed three types of movements derived from standard post-stroke rehabilitation protocols: (i) passive lifting, consisting of repeated vertical up-and-down motions of the arm; (ii) out-of-shoulder movements, involving arm motions that include shoulder rotation; and (iii) an activity-of-daily-living (ADL) task, in which subjects mimicked a feeding motion by moving the hand toward the mouth.
Each therapy type was performed under three interaction conditions designed to elicit different levels of subject participation: 1) passive condition: subjects were instructed to relax their arm and allow the robot to move it without resistance; 2) cooperative condition: subjects observed the motion demonstrated by the teleoperator wearing the AVP and actively attempted to minimize the interaction forces between their arm and the robot hand; 3) mixed condition: subjects alternated between relaxing their arm and actively cooperating with the robot motion. Together, these conditions were chosen to capture varying levels of subject participation and assistance, providing diverse interaction behaviors for learning contact-consistent impedance adaptation.
The study was approved by MIT’s Committee on the Use of Humans as Experimental Subjects (COUHES), with informed consent obtained from all participants. This dataset was used to train the model presented in this paper; however, the deployment of upper-limb rehabilitation applications is part of a separate project and is not reported here.
III-D Model Selection to Reconstruct sZFT
The model selection was guided by two initial requirements: (i) sufficient accuracy in reconstructing both translational and rotational components and (ii) a small model capacity, suitable for real-time deployment. To evaluate this trade-off, we varied hidden dimensions, number of layers, attention heads, and loss-weighting strategies.
Performance was assessed using translational error (positional loss in mm) and rotational error (angular magnitude $\theta$ about unit-axis and unit-axis deviation $\alpha$ in degrees).
@@ III-D Model Selection to Reconstruct sZFT
The initial requirements were $2\text{\,}\mathrm{mm}$ for the translational error and $2.8^{\circ}$ (half of the small-angle approximation) for the angular error and unit-axis deviation. These values were determined heuristically based on previous contact-rich experiments [37].
@@
Hidden Dim Heads Layers $\bm{\theta}$ Loss [deg] $\bm{\alpha}$ Loss [deg] Positional Loss [mm]
@@
A notable result was that retraining this model on a combined dataset of parkour and robotic-assisted therapy motions–two completely different physical interaction tasks–further improved accuracy (positional error $0.883$ mm, $\theta$ error $0.233^{\circ}$, $\alpha$ error $0.002^{\circ}$). Importantly, the two datasets were collected with different tools and inertial properties (lightweight tool for parkour and a 600 g prosthetic hand for physical therapy) yet the model compensated for these differences without explicit training. This robustness highlights the ability of Diffusion-Based Impedance Learning to generalize across diverse embodiements. The selected architecture was ultimately adopted as the final model because it satisfied the accuracy requirements while remaining small enough for real-time torque control.
Another notable finding of this work is the small amount of training data required. The model already showed strong performance on the parkour dataset with only tens of thousands of samples (17,00017{,}000 samples). For the combined data set, 72,00072{,}000 samples at a controller rate of 0.005s0.005\ $\text${s} correspond to just 1.6 hours of demonstration. This is in contrast to image-based Diffusion Models, which require 100,000100{,}000 samples even for learning single manipulation skills such as block pushing [10, 8].
III-E Energy-based Stiffness Estimation
Impedance control can be interpreted as virtual springs that pull the robot end-effector toward an equilibrium pose. When the robot interacts with the environment, the deformation of these virtual springs is transformed into external wrenches, which perform mechanical work. A shift of the reconstructed sZFT by the Diffusion Model changes the deformation, and thus the elastic energy stored in the virtual springs. By relating this change in elastic energy to the measured work, we can compute the stiffness equivalent that reproduces the observed interaction behavior.
For deployment, a user specifies baseline translational and rotational stiffness values that define the nominal interaction behavior. During execution, the Diffusion Model predicts the sZFT, which represents the equilibrium pose correction, consistent with the current interaction wrench. Rather than modifying the nominal ZFT, we adapt the impedance by reducing stiffness equivalent to the equilibrium reconstruction via sZFT.
Specifically, we compute nonnegative stiffness reduction terms $k_{t,i}^{\star}$ and $k_{r,i}^{\star}$ for each translational and rotational axis $i\in\{x,y,z\}$. These terms are subtracted from the predefined baseline stiffness values $K_{t,i,\text{max}}$ and $K_{r,i,\text{max}}$, yielding the effective stiffness used by the controller:
@@ III-E Energy-based Stiffness Estimation
In our experiments, the baseline stiffness values were set deliberately high ($K_{t,i,\text{max}}=800\ \text{N/m}$, $K_{r,i,\text{max}}=150\ \text{Nm/rad}$). High baseline stiffness is desirable for tasks that require significant elastic energy storage, such as pressing or inserting a workpiece into another. We therefore selected baseline stiffness values that the robot can realistically render and that are sufficient for all tasks considered. The Diffusion Model then autonomously reduced these stiffnesses when interaction occurred, adapting the impedance only where necessary.
III-E1 Translations
For translations, the displacement was defined as $\hat{\bm{e}}_{t}=\bm{p}-\hat{\bm{p}}_{0}$. The elastic energy $E_{t}\in\mathbb{R}$ of a virtual translational spring is
$E_{t}^{\text{spring}}=\tfrac{1}{2}\bm{K}_{t}\lVert\hat{\bm{e}}_{t}\rVert^{2}.$ (24)
Here, $\lVert.\lVert$ denotes the Euclidean norm. The work done by the external force $\bm{f}_{\text{ext}}\in\mathbb{R}^{3}$ along the virtual displacement $\hat{\bm{e}}_{t}$ is given by the inner product $\langle.,.\rangle$:
$E_{t}^{\text{work}}=\langle\bm{f}_{\text{ext}},\hat{\bm{e}}_{t}\rangle.$ (25)
Equating the right hand sites of Equation (24) with Equation (25) yields the translational stiffness contribution
$\bm{k}_{t}^{\star}=\frac{2\langle\bm{f}_{\text{ext}},\hat{\bm{e}}_{t}\rangle}{\lVert\hat{\bm{e}}_{t}\rVert^{2}}.$ (26)
To avoid oscillatory behavior, a damping term proportional to end-effector velocity was introduced:
$\tilde{\bm{e}}_{t}=\kappa_{t}\ \hat{\bm{e}}_{t}-\gamma_{t}\ \dot{\bm{p}},$ (27)
where $\kappa_{t}$ is a unit translational stiffness, $\dot{\bm{p}}$ is the translational end-effector velocity, and $\gamma_{t}$ a damping coefficient.
III-E2 Rotations
For rotations, the unit axis–angle representation $\hat{\bm{e}}_{r}=\hat{\bm{u}}_{0}\ \hat{\theta}_{0}$ was used (Equation (5)). The corresponding elastic energy $E_{r}\in\mathbb{R}$ of the virtual rotational spring is
$E_{r}^{\text{spring}}=\tfrac{1}{2}\bm{K}_{r}\lVert\hat{\bm{e}}_{r}\rVert^{2}.$ (28)
The work done by the external moment $\bm{m}_{\text{ext}}\in\mathbb{R}^{3}$ about the rotational displacement $\hat{\bm{e}}_{r}$ can be calculated by
$E_{r}^{\text{work}}=\langle\bm{m}_{\text{ext}},\hat{\bm{e}}_{r}\rangle.$ (29)
Equating the right hand sites of Equation (28) with Equation (29) results in the rotational stiffness contribution
$\bm{k}_{r}^{\star}=\frac{2\langle\bm{m}_{\text{ext}},\hat{\bm{e}}_{r}\rangle}{\lVert\hat{\bm{e}}_{r}\rVert^{2}}.$ (30)
Dynamic effects were addressed by augmenting the error with a damping term:
$\tilde{\bm{e}}_{r}=\kappa_{r}\ \hat{\bm{e}}_{r}-\gamma_{r}\ \bm{\omega},$ (31)
where $\kappa_{r}$ is a unit rotational stiffness, $\bm{\omega}$ is the angular end-effector velocity, and $\gamma_{r}$ a damping coefficient.
III-E3 Algorithm